# PoeTone: A Framework for Constrained Generation of Structured Chinese Songci with LLMs

### v1

@@ Abstract
This paper presents a systematic investigation into the constrained generation capabilities of large language models (LLMs) in producing Songci, a classical Chinese poetry form characterized by strict structural, tonal, and rhyme constraints defined by Cipai templates. We first develop a comprehensive, multi-faceted evaluation framework that includes: (i) a formal conformity score, (ii) automated quality assessment using LLMs, (iii) human evaluation, and (iv) classification-based probing tasks. Using this framework, we evaluate the generative performance of 18 LLMs, including 3 proprietary models and 15 open-source models across 4 families, under five prompting strategies: zero-shot, one-shot, completion-based, instruction-tuned, and chain-of-thought. Finally, we propose a Generate-Critic architecture in which the evaluation framework functions as an automated critic. Leveraging the critic’s feedback as a reward signal, we fine-tune 3 lightweight open-source LLMs via supervised fine-tuning (SFT), resulting in improvements of up to 5.88% in formal conformity. Our findings offer new insights into the generative strengths and limitations of LLMs in producing culturally significant and formally constrained literary texts.
@@ Introduction
This paper focuses on Songci, a prominent form of Chinese lyric poetry from the Song dynasty (960–1279 CE). Unlike Tang poetry, Songci follows fixed Cipai, predefined tune patterns that govern stanza count, line count, line length, tonal flow (píng and zè), rhyme type, and rhyme positions (Ge 2009). Effective Songci must not only conform to these structural constraints but also express coherent and emotionally resonant themes, typically involving longing, solitude, patriotism, or nature. This blend of metrical rigidity and artistic depth makes Songci a compelling testbed for constrained text generation. A common failure example in generating Songci is shown in Figure 1.
@@
[…]
@@
Related Word
Automatic generation of Chinese poetry has evolved from statistical and rule-based methods to neural networks (Wang et al. 2016; Chen and Cao 2024). Initial approaches used RNNs and LSTMs (Zhang and Lapata 2014; Wang et al. 2016), later enhanced by attention mechanisms to improve coherence and context modeling (Yi, Li, and Sun 2017).
A major research focus has been controlling poetic form and content. To handle strict structural and rhyming constraints, methods include dual-encoder models conditioned on rhythmic patterns (Luo et al. 2021) and form-aware generation guided by stressed weighting (Hu and Sun 2020). Thematic coherence has been addressed through planning-based structures (Wang et al. 2016), working memory models (Yi et al. 2018), salient clue guidance (Yi, Li, and Sun 2018; Gao et al. 2021), adversarial training for title consistency (Li et al. 2018), and graph neural networks for topic modeling (Yan et al. 2023). Sentiment and style control have been explored via multi-tag classification and conditional variational autoencoders (Shao et al. 2021; Chen et al. 2019; Song et al. 2025).
The rise of large language models (LLMs) has transformed Chinese poetry generation. An early study showed that a simple Generative Pre-trained Transformer (GPT) could produce high-quality classical poetry without extensive feature engineering (Liao et al. 2019). In further studies, researchers have fine-tuned pre-trained models on specialized poetry corpora, finding that smaller, domain-specific models can sometimes outperform larger, general-purpose LLMs like GPT-4o on poetic tasks (Tang et al. 2025; Wang, Guan, and Liu 2022). Innovations like CharPoet introduce token-free, character-level control to improve format accuracy (Yu et al. 2024), reflecting a trend to harness LLM capabilities while imposing classical poetic constraints.
Beyond text generation, interactive systems such as Jiuge (Zhipeng et al. 2019) and Yu Sheng (Ma, Zhan, and Wong 2023) incorporate human-in-the-loop frameworks for collaborative poem refinement. SongSong (Hu et al. 2025) generates music from Songci lyrics, addressing musical restoration, but leaving high-quality, metrically-correct Songci generation as an open challenge. Our work focuses on this core task, leveraging LLMs to generate formal Songci text.
@@
Each Cipai is encoded in machine-readable JSON format, capturing the following elements (illustrated in Figure 1):
@@
This dataset serves as a reference and training data for some of our analyses.
@@
We include 14 open-source LLMs from 4 major families:
@@
Ernie 4.5 Turbo zero-shot 4.75 4.05 4.10 2.15 3.45 3.10 3.53
@@
Model Cipai ID Topic Class. Source Attr.
@@
Figure 3: Expected formal improvements in generated Songci before and after fine-tuning. Red marks show errors; green marks show corrections.
@@
The iterable process proceeds as follows:
@@
Model Zero-shot One-shot Completion Instruction CoT
Llama-3.1-8B 36.83 (+1.63) 55.53 (+1.11) 54.77 (+1.37) 39.93 (+1.21) 41.30 (+1.17)
Mistral-Small 39.41 (+2.30) 57.81 (+3.40) 55.31 (+1.89) 38.96 (–1.46) 34.13 (–1.13)
Qwen3-8B 65.89 (+5.70) 70.41 (+5.88) 69.38 (+5.63) 54.31 (+0.33) 59.47 (–3.11)

### v2

@@ Abstract
This paper presents a systematic investigation into the constrained generation capabilities of large language models (LLMs) in producing Songci, a classical Chinese poetry form characterized by strict structural, tonal, and rhyme constraints defined by Cipai templates. We first develop a comprehensive, multi-faceted evaluation framework that includes: (i) a formal conformity score, (ii) automated quality assessment using LLMs, (iii) human evaluation, and (iv) classification-based probing tasks. Using this framework, we evaluate the generative performance of 18 LLMs, including 3 proprietary models and 15 open-source models across 4 families, under five prompting strategies: zero-shot, one-shot, completion-based, instruction-based, and chain-of-thought. Finally, we propose a Generate-Critic architecture in which the evaluation framework functions as an automated critic. Leveraging the critic’s feedback as a scoring function for best-of-N selection, we fine-tune 3 lightweight open-source LLMs via supervised fine-tuning (SFT), resulting in improvements of up to 5.88% in formal conformity. Our findings offer new insights into the generative strengths and limitations of LLMs in producing culturally significant and formally constrained literary texts.
Codehttps://github.com/ZhanQu945/PoeTone
@@ Introduction
This paper focuses on Songci, a prominent form of Chinese lyric poetry from the Song dynasty (960–1279 CE). Unlike Tang poetry, Songci follows fixed Cipai, predefined tune patterns that govern stanza count, line count, line length, tonal flow (píng and zè), rhyme type, and rhyme positions (Ge 2009). Effective Songci must not only conform to these structural constraints but also express coherent and emotionally resonant themes, typically involving longing, solitude, patriotism, or nature. This blend of metrical rigidity and artistic depth makes Songci a compelling testbed for constrained text generation. A common failure case in generating Songci is shown in Figure 1.
@@
Third, we propose a Generate-Critic architecture that uses rule-based conformity scoring to guide fine-tuning. Applied to 3 open-source LLMs via best-of-$N$ rejection sampling and LoRA-based supervised fine-tuning (SFT), this method improves adherence to structural and tonal constraints by up to 5.88%.
@@
Related Work
Automatic generation of Chinese poetry has evolved from statistical and rule-based methods to neural networks (Wang et al. 2016; Chen and Cao 2024). Initial approaches used RNNs and LSTMs (Zhang and Lapata 2014; Wang et al. 2016), later enhanced by attention mechanisms to improve coherence and context modeling (Yi et al. 2017).
A major research focus has been controlling poetic form and content. To handle strict structural and rhyming constraints, methods include dual-encoder models conditioned on rhythmic patterns (Luo et al. 2021) and form-aware generation guided by stressed weighting (Hu and Sun 2020). Thematic coherence has been addressed through planning-based structures (Wang et al. 2016), working memory models (Yi et al. 2018b), salient clue guidance (Yi et al. 2018a; Gao et al. 2021), adversarial training for title consistency (Li et al. 2018), and graph neural networks for topic modeling (Yan et al. 2023). Sentiment and style control have been explored via multi-tag classification and conditional variational autoencoders (Shao et al. 2021; Chen et al. 2019; Song et al. 2025).
The rise of large language models (LLMs) has transformed Chinese poetry generation. An early study showed that a simple Generative Pre-trained Transformer (GPT) could produce high-quality classical poetry without extensive feature engineering (Liao et al. 2019). In further studies, researchers have fine-tuned pre-trained models on specialized poetry corpora, finding that smaller, domain-specific models can sometimes outperform larger, general-purpose LLMs like GPT-4o on poetic tasks (Tang et al. 2025; Wang et al. 2022). Innovations like CharPoet introduce token-free, character-level control to improve format accuracy (Yu et al. 2024), reflecting a trend to harness LLM capabilities while imposing classical poetic constraints.
Beyond text generation, interactive systems such as Jiuge (Zhipeng et al. 2019) and Yu Sheng (Ma et al. 2023) incorporate human-in-the-loop frameworks for collaborative poem refinement. SongSong (Hu et al. 2025) generates music from Songci lyrics, addressing musical restoration, but leaving high-quality, metrically-correct Songci generation as an open challenge. Our work focuses on this core task, leveraging LLMs to generate formal Songci text.
@@
Each Cipai is encoded in machine-readable JSON format, capturing the following elements:
@@
This dataset serves as reference and training data for some of our analyses.
@@
We include 15 open-source LLMs from 4 major families:
@@
ERNIE 4.5 Turbo completion 4.75 4.05 4.10 2.15 3.45 3.10 3.53
@@
Model Cipai ID Theme Class. Source Attr.
@@
Figure 3: Expected formal improvements in generated Songci before and after fine-tuning.
@@
The iterative process proceeds as follows:
@@
Llama-3.1-8B Mistral-Small Qwen3-8B
Zero-shot 36.83 (+1.63) 39.41 (+2.30) 65.89 (+5.70)
One-shot 55.53 (+1.11) 57.81 (+3.40) 70.41 (+5.88)
Completion 54.77 (+1.37) 55.31 (+1.89) 69.38 (+5.63)
Instruction 39.93 (+1.21) 38.96 (–1.46) 54.31 (+0.33)
CoT 41.30 (+1.17) 34.13 (–1.13) 59.47 (–3.11)
@@ Conclusion
Acknowledgments
The authors acknowledge the financial support by the Federal Ministry of Research, Technology and Space of Germany and by Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus in the programme Center of Excellence for AI-research “Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig”, project identification number: ScaDS.AI.
Additionally, the project was supported by the German Federal Ministry of Research, Technology and Space (BMFTR) via the Software Campus project (01—S23070).
The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Research, Technology and Space of Germany and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner).